OpenStudy (anonymous):

password consists of 2 letters followed by 2 digits. How many different passwords can be formed?

jimthompson5910 (jim_thompson5910):

There are 26 letters If repeats are allowed, then there are 26*26 = 676 ways to pick a 2 letter combo There are 10 numeric digits Again if repeats are allowed, then there are 10*10 = 100 ways to pick a 2 digit number combo ---------------------- Multiply the two results: 676*100 = 67600 So there are 67600 different passwords possible

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (whpalmer4):

this of course assumes that "A" and "a" are not treated differently. If they are (and passwords are usually case sensitive these days) then you have 52*52 = 2704 2 letter combinations, or 2704*100=270400 possible different passwords

jimthompson5910 (jim_thompson5910):

true, didn't think about case sensitive passwords

OpenStudy (whpalmer4):

quite possible the problem author didn't, either :-)

jimthompson5910 (jim_thompson5910):

good point, they probably just want to include one copy of each letter

OpenStudy (whpalmer4):

and we're assuming English, not say Swedish :-)